Author Topic: [SOLVED] Datron Wavetek 4708 AC problem / Fail 2,5,7  (Read 854 times)

0 Members and 1 Guest are viewing this topic.

Offline doktor pyta

  • Frequent Contributor
  • **
  • Posts: 377
  • Country: pl
    • RF Scientific
[SOLVED] Datron Wavetek 4708 AC problem / Fail 2,5,7
« on: May 04, 2019, 04:21:30 pm »
I wonder if someone had similar problem before.
After switching on the calibrator works OK.
After say 15min warmup and using HV AC ranges fails start to appear.
They are present ONLY on 10kHz and 100kHz frequency ranges. So maybe it is related to HF transformer.

When I type 700V 10kHz and press 'output on' sometimes the correct voltage is present and more often I see Fail 2 (for only 0.2s) then Fail 5 and after prressing Reset : Fail 7.
First I thought that it is due to some failure in the output protection circuits. I checked them and even swapped main IC with the new ones. It didn't help.
Next bright thought was that maybe there is something that triggers these protection circuits...
I connected oscilloscope with tek P6015 HV probe to the output and I really saw some spikes exceeding normal levels.

Sometimes they occur at turn on and nothing happpens- output is present.
Sometimes at turn on they trigger Fail sequence mentioned above.
Sometimes they occur during pressing 'output off' and they do nothing or trigger 'Error OL' message.

The spikes look like oscillations at frequency in 100kHz range.
Below some screenshots.

I would greatly appreciate any suggestions from You.
« Last Edit: November 27, 2020, 10:11:40 am by doktor pyta »
 

Offline doktor pyta

  • Frequent Contributor
  • **
  • Posts: 377
  • Country: pl
    • RF Scientific
Re: [SOLVED] Datron Wavetek 4708 AC problem / Fail 2,5,7
« Reply #1 on: November 27, 2020, 10:24:58 am »
That was long ago, but maybe my comment will help someone in the future.

There were two issues.

1. dead LED section of M3 optocoupler on REF Divider board.
It carries 16kHz reference clock for PLL in AC synthesizer.
When this signal is not present, the amplitude and frequency of AC are behaving very strange.
I think it was damaged by capacitively coupled spikes which occurred during power disconnecting in the moments the instrument detected Fail.
I put 1N4148  antiparallel to the LED to avoid this in the future as it happened two times !

2. Bias adj. pot R10 on Positive heatsink assy was not set properly. Approx. 3 turns from the place it should be.
This caused crossover distortion when PA was not loaded (that's how I tracked it) and caused HF oscillations when HF transformer was connected to the output of the PA.
To correct that I just followed Bias adjustment procedure described in the service manual.

Update: All problems with bias current was due to leaky tant C7 on positive heatsink. The only one capacitor I forgot to replace...

Cheers!
« Last Edit: November 27, 2020, 07:52:22 pm by doktor pyta »
 
The following users thanked this post: ManateeMafia, volvo_nut_v70, LaserEng, TUMEMBER, syau


Share me

Digg  Facebook  SlashDot  Delicious  Technorati  Twitter  Google  Yahoo
Smf